OpenStudy (anonymous):

x^2 -y^2/xy /1/y - 1/x=

OpenStudy (anonymous):

\[\Large \frac{\frac{x^2-y^2}{xy}}{\frac{1}{y}-\frac{1}{x}}\]This?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, so I'd start with simplifying: \[\frac{1}{y}-\frac{1}{x}\]

OpenStudy (anonymous):

Do you know how to do this?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

In order to subtract two fractions, they have to have a common denominator.

OpenStudy (anonymous):

so is it 1

OpenStudy (anonymous):

\[\frac{1}{y}-\frac{1}{x} = \frac{1}{y} \cdot \frac{x}{x} -\frac{1}{x} \cdot \frac{y}{y} = \frac{x}{xy} - \frac{y}{xy} = \frac{x-y}{xy}\]

OpenStudy (anonymous):

Does what I did make sense to you?

OpenStudy (anonymous):

We end up with: \[\Large \frac{\frac{x^2-y^2}{xy}}{\frac{x-y}{xy}}\]

OpenStudy (anonymous):

yes I was looking it over. So i have to pair them up inoder to get and answer... Sorry my son had just hit his head on the wall that is why a delay in responding back.

OpenStudy (anonymous):

Do you know what to do from here?

OpenStudy (anonymous):

I subtract and then divide it by xy then do the same for the bottom to get my answer

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

xy^2

OpenStudy (anonymous):

That is not correct, how did you get that?

OpenStudy (anonymous):

I thought it was simplifying..

OpenStudy (anonymous):

to me i thought it was one the because both the top and the bottom equal out to me

OpenStudy (anonymous):

Dividing by a fraction is the same as multiplying by its reciprocal.\[ \Large \frac{\frac{x^2-y^2}{xy}}{\frac{x-y}{xy}} = \frac{x^2-y^2}{xy}\cdot \frac{xy}{x-y} \]

OpenStudy (anonymous):

In this case the xy cancel out leaving: \[\frac{x^2-y^2}{x-y}\]

OpenStudy (anonymous):

This can also be simplified more though, so it's not done.

OpenStudy (anonymous):

We have to factor \(x^2-y^2\) to see if it has \(x-y\) as a factor.